Find the inverse of the function.
f(x)=4x^3-1

Answer by edjones(7311)

f(x)=4x^3-1
y=4x^3-1 sustitute y for f(x)
x=4y^3-1 Switch x and y
4y^3=x+1 solve for y
y^3=(x+1)/4
y=[(x+1)/4]^(1/3) Take the cube root of each side.
f^-1(x)=[(x+1)/4]^(1/3)
.
Ed